Large Manufacturing, Inc. is considering investing in some new equipment whose data are shown below. The equipment has a 3-year class life and will be depreciated by the MACRS depreciation system, and it will have a positive pre-tax salvage value at the end of Year 3, when the project will be closed down. Also, some new working capital will be required, but it will be recovered at the end of the project's life. Revenues and cash operating costs are expected to be constant over the project's 3-year life. What is the project's Initial Cash Outlay at time 0? I know this answer, its 80000. Need the below questions answered.

WACC

11.0%

Net investment in fixed assets (depreciable basis)

$70,000

Required new working capital

$10,000

Sales revenues, each year

$95,000

Cash operating costs excl. depr'n, each year

$30,000

Expected pretax salvage value

$9,000

Tax rate

30.0%

Initial Cash Outlay at time 0 = 80000

what is the Year 1 Net Operating Cash Flow?  

what is the Terminal Year Non–Operating Cash Flow at the end of Year 3?

Thank you!
Answer #1
1] Net investment in fixed assets $           70,000
+New working capital $           10,000
=Initial cash outlay at time 0 $           80,000
2] 0 1 2 3 4 Total
Annual sales revenue $            95,000 $             95,000 $           95,000
-Cash operating costs [excluding depreciation] $            30,000 $             30,000 $           30,000
-Depreciation $            23,331 $             31,115 $           10,367 $          5,187 $        70,000
=Net operating income $            41,669 $             33,885 $           54,633
-Tax at 30% $            12,501 $             10,166 $           16,390
=NOPAT $            29,168 $             23,720 $           38,243
+Depreciation $            23,331 $             31,115 $           10,367
=Operating cash flow $            52,499 $             54,835 $           48,610
3] Pretax salvage value $             9,000
Book value at end of year 3 $             5,187
Gain on sale $             3,813
Tax on sale at 30% $             1,144
After tax salvage value = 9000 -1144 = $             7,856
Realization of working capital $           10,000
Terminal non operating cash flow at end of year 3 = 7856+10000 = $           17,856
4] Annual project cash flows $         -80,000 $            52,499 $             54,835 $           66,466
PVIF at 11% [PVIF = 1/1.11^t] 1 0.90090 0.81162 0.73119
PV at 11% $         -80,000 $            47,297 $             44,505 $           48,600
NPV $           60,401
5] As the NPV is positive, the project can be implemented.